A sample of oxygen gas has a volume of 2.1 L at 3.6 atm pressure and 200 K temperature.

What would the volume of this sample be at 1.0 atm and 400 K?


How many liters ?


1
Expert's answer
2021-10-20T02:53:08-0400

"V_1 = 2.1 \;L \\

p_1 = 3.6 \;atm \\

T_1=200 \;K \\

p_2 = 1.0 \;atm \\

T_2 = 400 \;K \\

\frac{p_1V_1}{T_1} = \frac{p_2V_2}{T_2} \\

V_2 = \frac{p_1V_1T_2}{T_1p_2} \\

V_2 = \frac{3.6 \times 2.1 \times 400}{200 \times 1.0} = 15.12 \;L"

Answer: 15.12 L